OpenStudy (anonymous):

Given the linear equation y=3/4x-2, find the y-coordinates of the points (-8, ), (-4, ), and (4, ). Please show all of your work. Plot those points and graph the linear equation.

OpenStudy (helder_edwin):

just plug into the formula the given values of x. for instance: x=4 then \[ \large y=\frac{3}{4}x-2=\frac{3}{4}4-2=3-2=1 \] then (4,1)